# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ V-C Stability (RQ3)
Results presented in the preceding chapters demonstrate that SeSE effectively detects hallucinations in the question-answering process of various LLMs and exhibits strong generalization across multiple domains. An additional crucial aspect is the stability of SeSE’s performance as a hallucination detector for LLMs. Given the inherent stochasticity of LLM generation [49], it is essential to evaluate whether SeSE’s performance is influenced by these uncertainties. To this end, we conduct five repeated experiments with distinct random seeds across 25 experimental scenarios (25 model-dataset pairs), resulting in a total of 125 runs. For each run, a 95% confidence interval is computed using 1,000 bootstrap resampling. While standard errors are reported in Table I, it should be noted that, in the context of LLM uncertainty quantification, standard errors tend to depend more on the LLM and dataset than on the UQ method itself. Rather than absolute values, the consistency of relative results across scenarios serves as a more reliable indicator of performance variation [16]. Therefore, we adopt the pairwise win rate as the primary evaluation metric and assess statistical significance using a binomial test. Within each scenario, the method with more wins across the five repeated trials is deemed superior. The heatmaps in Figure 5 visualize the pairwise win rates, showing that SeSE consistently outperforms baselines at a significance level of $p<0.05$. This finding indicates that, although SeSE’s performance may vary across scenarios, its comparative advantage remains highly consistent, making it a more stable uncertainty estimator in practical applications.
V-D Sensitivity to Hyperparameters (RQ4)
V-D1 Sensitivity to Sampling Numbers
The number of sampled responses $N$ does not need to be large. Figure 6(a) illustrates how the hallucination detection performance varies with $N$ for sentence-length generations, while Figure 6(b) presents the corresponding results for long-form generations. The reported values are aggregated across all datasets and LLMs. For sentence-length generations, performance gains show diminishing returns at $N\approx 5$. However, increasing $N$ to 10 can still be beneficial. For long-form generation, we find that optimal performance is achieved with $9\text{--}10$ responses. Different from sentence-level generation, more generations don’t always improve performance. This occurs because the stochastic decoding strategies are commonly employed to boost the output diversity of LLMs [49], which increases the likelihood of selecting low-probability tokens. Excessively increasing sample numbers amplifies the selection probability of low-probability tokens, potentially introducing irrelevant content and thereby diminishing the relative weight of original claims from greedy decoding. In this study, we set $N=10$ for both sentence-length and long-form experiments.
(a) Number of sentence-length generations used for entropy.
@@ V-D1 Sensitivity to Sampling Numbers
Figure 6: RQ4.1: The performance of SeSE with different number of sampled responses (NN).
@@
Figure 7: RQ4.2: The performance of SeSE when adopting different tree height $K$.
V-D2 Sensitivity to Encoding Tree Height
@@
Figure 7 shows SeSE’s performance sensitivity to encoding tree height $K$. The accuracy improvement of the best tree height compared with $K=1$ is annotated in the figure. When $K=1$, SeSE degenerates to the non-hierarchical von Neumann graph entropy [31]. Compared with $K=1$, SeSE achieves optimal performance on two and three datasets at $K=2$ and $K=3$, respectively. We also observe that the optimal $K$ correlates with task difficulty. For instance, on the simpler dataset TriviaQA, SeSE peaks at $K=2$, while on the challenging SQuAD, the optimum is found at $K=3$. These findings demonstrate that SeSE can flexibly adapt to diverse downstream tasks with optional encoding tree depth. We provide unaggregated results for Figure 6 and additional statistics with detailed analyses for Figure 7 in Appendix VIII-D4.
V-E Ablation Study
Table III presents the ablation study on hallucination detection performance under different graph uncertainty metrics and semantic graph construction schemes. Specifically, in the experiments on graph uncertainty metrics, only the structural entropy is replaced with alternative metrics, including Eigenvalue, Eccentricity, Spectral Gap and Degree [19]. SeSE-undir denotes the use of an undirected semantic graph, and SeSE-com refers to a complete semantic graph. Among all variants, the complete SeSE achieves the best results across all experimental settings. First, SeSE outperforms all other graph uncertainty metrics. They essentially simplify the graph into a collection of nodes and edges with specific probabilistic distributions, overlooking critical semantic topological structures that convey the core graph information. In contrast, SeSE formalizes hierarchical abstraction through an encoding tree, effectively capturing the semantic structural information. Second, removing either directed semantic modeling or graph sparsification results in a decline in detection performance. As shown in Figure 8, the undirected complete graph (see (a)) contains numerous edges with near-zero weights and obfuscate the asymmetric nature of semantic spaces (see (b)). Only the“dir” + “sparsification” combination (see (c)) optimally models the semantic space, effectively revealing both core semantic distributions and directional relationships. Detailed definitions of the used graph uncertainty metrics are provided in Appendix Table VIII.
TABLE III: Ablation study of SeSE under different setups.
Llama-3.2-3B Llama-3.1-8B Llama-3.1-70B Qwen-3-4B Qwen-3-30B
AUROC AURAC AUROC AURAC AUROC AURAC AUROC AURAC AUROC AURAC
Eccentricity 0.6664 0.5059 0.6305 0.4989 0.7259 0.7343 0.6919 0.5707 0.6773 0.6547
Spectral Gap 0.7187 0.5445 0.6530 0.5173 0.7705 0.7549 0.7546 0.6090 0.7479 0.6924
Eigenvalue 0.7833 0.5670 0.7007 0.5314 0.8027 0.7675 0.7913 0.6216 0.7785 0.6992
Degree 0.8117 0.5665 0.7247 0.5342 0.8211 0.7639 0.7979 0.6286 0.7953 0.6892
SeSE 0.8145 0.5783 0.7356 0.5495 0.8330 0.7775 0.8266 0.6349 0.7992 0.7103
SeSE-undir 0.7711 0.5585 0.7067 0.5250 0.8069 0.7514 0.7925 0.6127 0.7720 0.6872
SeSE-comple 0.7193 0.5220 0.6946 0.5177 0.7462 0.7110 0.7855 0.6098 0.7641 0.6841
  • The best results are bolded.
Figure 8: The adjacent matrix of the semantic space with different graph modeling methods. A gray grid represents that there is no edge between response $i$ and $j$.
V-F Case Study
We present two representative examples from the SQuAD dataset in Figure 9. To analyze cases where SE and SeSE are inconsistent, we ranked all cases by SE and selected two examples with similar indices but different factual labels (Hallu/Non-hallu). As shown in Figure 9(a), when semantic structures are ignored, the semantic spaces of these two cases appear superficially similar, both divisible into five semantic clusters with similar sizes. However, since SE solely estimates the Shannon entropy of semantic cluster probability distributions, it underestimates the divergence between these two cases, yielding a value of only 0.14. In contrast, SeSE employs an optimal $K$-dimensional encoding tree $\mathcal{T}^{*}$ constructed according to the principle of structural information minimization, which captures the intrinsic hierarchical organization of the semantic space, i.e., compressing it most succinctly. The structural entropy of $\mathcal{T}^{*}$ quantifies the minimal number of bits required to describe the semantic space [31]. As illustrated in Figures 9(a)–(c), increasing the depth of the encoding tree reveals that the hallucination case exhibits greater resistance to compression due to its disordered semantic structure, resulting in a final compression ratio of only 36%. Conversely, the non-hallucinated case, characterized by more regular semantic organization, is more readily compressed, achieving a final ratio of 46%. This difference increases the entropy gap $\Delta_{SeSE}$ from 0.25 to 0.45. Consequently, SeSE effectively identifies fine-grained uncertainty distinctions even when existing semantic UQ methods fail [16, 17, 19]. Detailed information of the used cases is provided in Appendix VIII-E.
Figure 9: Two cases in which SeSE effectively identifies subtle uncertainty when SE fails. $\Delta_{SeSE}$ represents the SeSE difference between the encoding tree of the $1^{st}$ and $2^{nd}$ row. $\Delta_{\downarrow}^{\%}$ represents the compression ratio compared to the 1-dimensional encoding tree (i.e., part (a)).
VI Related Work
VI-A Uncertainty Estimation in LLMs
This section reviews the fundamental ideas of existing UQ methods and their potential limitations, which are then addressed in the development of SeSE.
@@ VI-A Uncertainty Estimation in LLMs
The first direction is to conduct supervised learning, either by fine-tuning the base LLMs or by adding additional layers to obtain uncertainty scores [50, 51]. Despite their promise, as these methods are often model-specific and require additional training, they have poor generalization ability and cannot be applied to closed-source models. In contrast, SeSE offers an “off-the-shelf” solution applicable across both open- and closed-source models without requiring any architectural modifications or additional training. The second line of research seeks to prompt LLMs to express their uncertainty about their own responses using natural language [24, 52, 22]. However, as noted by Kalai et al. [53], since most existing evaluation methods fail to incentivize models to express uncertainty honestly, LLMs tend to exhibit high confidence even when they might provide incorrect outputs. Overall, the performance of verbalized confidence is generally inferior to other more advanced methods [22].
The aforementioned methods primarily focus on token-level uncertainty, neglecting the model’s semantic uncertainty. In fact, semantic uncertainty is crucial for analyzing the hallucinations in LLM outputs, as it directly influences the accuracy of the models’ outputs. As a third direction, semantic entropy [16, 14, 15] is one of the most advanced metrics for quantifying uncertainty in the semantic space. It groups responses into semantic clusters and calculates the Shannon entropy of the resulting cluster probability distribution as the uncertainty metric. Although its performance generally surpasses that of other token-based approaches, it considers only the one-cut semantic equivalence relationship. Subsequent works like kernel language entropy [17] tried to address this issue using semantic graph kernels, but still suffer from two inherent limitations: (1) the adopted kernel functions make it challenging to interpret the model’s uncertainty; (2) semantic analysis is conducted at the same level without hierarchy, which violates the principle of “compositional similarity” and limits the ability to distinguish between fine-grained uncertainties. Based on structural information theory, our proposed SeSE enables fine-grained semantic analysis through the hierarchical encoding tree, thereby providing more precise uncertainty estimates.
VI-B Structural Entropy
Structural entropy, originally introduced by Li and Pan [29], serves as a metric for graph uncertainty. It generalizes Shannon entropy from unstructured probability distributions to graphs with arbitrary topologies. Structural information theory employs an encoding tree to encode graphs, and minimizing structural entropy can reveal the intrinsic structure of graphs. In recent years, it has been widely employed in diverse domains, including graph kernels [54], text classification [32], social detection [55], and semi-supervised clustering [56]. However, most existing research on structural information focuses on undirected graphs, which fail to capture the irreversible logical directional relationships between responses, leading to inevitable information loss. In this study, we optimize the native structural entropy for directed graphs to facilitate hierarchical abstracting under asymmetric scenarios.
VII Conclusion
This paper proposes SeSE, a principled UQ framework based on structural information theory, capturing semantic structural information within semantic spaces to provide more precise uncertainty estimates. First, we develop an adaptively sparsified directed semantic graph construction algorithm to effectively model semantic spaces. Then, we perform hierarchical abstraction using an encoding tree to quantify the inherent semantic uncertainty of LLMs. Additionally, we extend SeSE to quantify claim-level uncertainty in long-form outputs, providing interpretable uncertainty estimates. SeSE outperforms existing competitors and is verified to be a reliable UQ framework for LLM hallucination detection. Notably, SeSE is an “off-the-shell”, opening up new possibilities for reliably deploying LLMs in resource-constrained environments.
Appendix
VIII-A Theoretical Proofs
In this section, we prove proposed Theorem 1 using the Perron-Frobenius theorem.
Lemma 1 (Perron–Frobenius Theorem).
Let $M\in\mathbb{R}^{n\times n}$ be an irreducible non-negative matrix. Then:
  1. (1)
    $\rho(M)=\max\{|\lambda|:\lambda\in\sigma(M)\}$ is an eigenvalue of $M$;
  2. (2)
    There exists a positive vector $\mathbf{x}>0$ such that $M\mathbf{x}=\rho(M)\mathbf{x}$;
  3. (3)
    The algebraic multiplicity of $\rho(M)$ is $1$;
  4. (4)
    $M^{\top}$ is also irreducible and non-negative.
Given that $A$ is a stochastic matrix, all its row sums equal $1$. Hence, the all-ones vector $\mathbf{e}=(1,1,\dots,1)^{\top}$ satisfies
$A\mathbf{e}=\mathbf{e},$ (18)
which implies that $1$ is an eigenvalue of $A$, i.e., $1\in\sigma(A)$. Therefore, the spectral radius $\rho(A)=\max\{|\lambda|:\lambda\in\sigma(A)\}$ satisfies
$\rho(A)\geq 1.$ (19)
For any matrix $M$, the spectral radius is bounded above by its $\infty$-norm: $\rho(M)\leq\|M\|_{\infty}$, where $\|M\|_{\infty}=\max_{i}\sum_{j}|M_{ij}|$. Since $A$ is stochastic, we have $\|A\|_{\infty}=\max_{i}\sum_{j}A_{ij}=1$, and thus
$\rho(A)\leq 1.$ (20)
Combining both inequalities yields $\rho(A)=1$.
A stationary distribution $\pi^{\top}$ is defined by the condition $\pi^{\top}A=\pi^{\top}$, which is equivalent to
$A^{\top}\pi=\pi.$ (21)
Because $A$ is irreducible and non-negative, its transpose $A^{\top}$ is also irreducible and non-negative. Applying the Perron–Frobenius Theorem to $A^{\top}$, we conclude:
  1. The spectral radius $\rho(A^{\top})=1$ is an eigenvalue of $A^{\top}$;
  2. There exists a strictly positive eigenvector $\pi>\mathbf{0}$ such that $A^{\top}\pi=\pi$;
  3. The eigenvalue $1$ has algebraic multiplicity one, so $\pi$ is unique up to scalar multiplication.
We now normalize $\pi$ by its $\ell_{1}$-norm:
$\hat{\pi}=\frac{\pi}{\pi^{\top}\mathbf{e}},$ (22)
where $\pi^{\top}\mathbf{e}=\sum_{i}\pi_{i}>0$ since $\pi>\mathbf{0}$. By construction, $\hat{\pi}^{\top}\mathbf{e}=\dfrac{\pi^{\top}\mathbf{e}}{\pi^{\top}\mathbf{e}}=1$, and
$\hat{\pi}^{\top}A=\left(\frac{\pi}{\pi^{\top}\mathbf{e}}\right)^{\top}A=\frac{\pi^{\top}A}{\pi^{\top}\mathbf{e}}=\frac{\pi^{\top}}{\pi^{\top}\mathbf{e}}=\hat{\pi}^{\top}.$
Hence, $\hat{\pi}^{\top}$ is a stationary distribution.
To prove uniqueness, suppose $\sigma^{\top}$ is another stationary distribution, i.e., $\sigma^{\top}A=\sigma^{\top}$, which implies $A^{\top}\sigma=\sigma$. Then $\sigma$ is a positive eigenvector of $A^{\top}$ corresponding to the eigenvalue $1$. Since the eigenspace is one-dimensional, $\sigma$ must be a scalar multiple of $\pi$. After normalization, we obtain $\sigma=\hat{\pi}$, confirming the uniqueness of $\hat{\pi}$.
Therefore, $G_{\text{dir}}$ admits a unique stationary distribution $\hat{\pi}^{\top}$, which is the unique normalized eigenvector of $A^{\top}$ corresponding to the dominant eigenvalue $1$. ∎
VIII-B Hierarchical Abstracting

### v3

@@ 4.2 Main Results
While standard errors are reported in Table 1, it should be noted that in the context of LLM UQ, standard errors tend to depend more on the LLM and the dataset than on the UQ method itself Farquhar et al. [2024]. Rather than absolute values, the consistency of relative results across scenarios serves as a more reliable indicator of performance variation Farquhar et al. [2024]. Therefore, we perform a binomial statistical significance test across all baselines. First, we conduct five repeated experiments with distinct random seeds across 20 experimental scenarios (100 runs). For each run, we compute the 95% confidence interval using 1,000 bootstrap resamples. We then adopt the pairwise win rate as the primary evaluation metric. Within each scenario, the method with more wins across five repeated trials is deemed superior. The heatmaps in Figure 2 visualize the pairwise win rates, showing that SeSE consistently outperforms baselines at a significance level of $p<0.05$. This finding indicates that although the performance of SeSE may vary across scenarios, its comparative advantage remains highly consistent, making it a more stable uncertainty estimator in practical applications.
(a) Number of short-form generations used for entropy.
@@
Figure 3: The performance of SeSE with different number of sampled responses (NN).
@@
Figure 4: The performance of SeSE when adopting different tree height $K$.
4.3 Ablation Experiments
Number of Sampling Numbers
The number of sampled responses $N$ does not need to be large. Figure 3 illustrates how the performance of uncertainty quantification varies with $N$ for both short-form and long-form generations. The reported values are aggregated across all datasets and LLMs. For short-form generations, performance gains show diminishing returns at $N\approx 5$. However, increasing $N$ to 10 can still be beneficial. For long-form generation, we find that optimal performance is achieved with $9\text{-}10$ responses. Different from short-form scenarios, more generations don’t always improve performance. This occurs because the stochastic decoding strategy of LLMs increases the likelihood of selecting low-probability tokens. Excessively increasing sample numbers amplifies the selection probability of selecting such tokens, potentially introducing irrelevant content and thereby diminishing the relative weight of original greedily-decoding claims. We provide unaggregated results for Figure 3 in Appendix D.7.
Encoding Tree Height
@@ 4.3 Ablation Experiments
Figure 4 shows SeSE’s performance sensitivity to the encoding tree height $K$. The accuracy improvement of the best tree height compared with $K=1$ is annotated in the figure. When $K=1$, SeSE degenerates to the non-hierarchical Shannon entropy of the graph’s stationary distribution. Compared with $K=1$, SeSE achieves its optimal performance on two and three datasets at $K=2$ and $K=3$, respectively. We also observe that the optimal $K$ correlates with task difficulty. For instance, on the simpler dataset TriviaQA, SeSE peaks at $K=2$, while on the challenging SQuAD, the optimum is found at $K=3$. These findings demonstrate that SeSE can flexibly adapt to diverse downstream tasks with optional encoding tree depth. Additional statistics and detailed analyses for Figure 4 are provided in Appendix D.7.
5 Related Work
Short-form Uncertainty Estimation in LLMs
@@ 5 Related Work
Recently, numerous UQ methods have emerged. A primary direction involves supervised learning, either by fine-tuning base LLMs or adding external layers to predict uncertainty scores Liu et al. [2024], Xie et al. [2024], Li et al. [2025a]. Despite their promise, these methods are typically model-specific and cannot be applied to closed-source models, resulting in limited scalability and availability. Another line of research explores verbalized uncertainty, prompting LLMs to express uncertainty via natural language Kadavath et al. [2022], Tian et al. [2023b], Mohri and Hashimoto [2024], Wang et al. [2025]. However, since most existing evaluation methods fail to incentivize models to express uncertainty honestly Kalai et al. [2025], LLMs tend to exhibit overconfidence even when generating incorrect outputs. Consequently, verbalized uncertainty underperforms probabilistic methods Mohri and Hashimoto [2024].
The aforementioned methods primarily focus on lexical uncertainty, neglecting semantic uncertainty, which is a more essential indicator of LLM trustworthiness as it directly reflects response correctness Farquhar et al. [2024]. Semantic Entropy Farquhar et al. [2024] represents a significant advance that calculates the Shannon entropy of semantic equivalence clusters as an uncertainty metric. However, SE is a binary, one-cut measurement that overlooks finer semantic differences between responses. While recent follow-ups like Kernel Language Entropy (KLE) Nikitin et al. [2024] and Semantic Graph Density (SGD) Xiao et al. [2025] model fine-grained semantic relationships in semantic graphs, they fail to capture the semantic topological structure which reflects the informational essence of graphs Li et al. [2024], Su et al. [2025], limiting their ability to distinguish between fine-grained uncertainties. To address this, SeSE quantifies the uncertainty inherent in the semantic graph after optimal compression based on the structural entropy minimization principle, providing more precise and interpretable uncertainty estimates.
Granular uncertainty estimation
Quantifying uncertainty in long-form generation has attracted increasing attention [Manakul et al., 2023, Mohri and Hashimoto, 2024, Zhang et al., 2024, Jiang et al., 2024]. Most relevant to our work is Graph Uncertainty [Jiang et al., 2024], which employs graph centrality metrics like degree and closeness as heuristic proxies for uncertainty. While SeSE also operates on a bipartite semantic graph in long-form settings, the key distinction is that SeSE exploits latent semantic structural dependencies from a random walk perspective, providing interpretable uncertainty estimates for black-box LLMs.
Structural Information Theory
Structural information theory Li and Pan [2016] provides a framework to quantify dynamic uncertainty in complex networks. Minimizing structural entropy, by searching for a nested partitioning tree, provides a theoretically grounded method to identify the intrinsic community structure of a network Li [2024]. This theory has been successfully applied across diverse domains, including graph learning Yang et al. [2025], social networks Yang et al. [2024], and reinforcement learning Zeng et al. [2024].
6 Conclusion
In this study, we proposed SeSE, a principled black-box UQ framework that works both for open- and closed-source LLMs. By constructing the optimal hierarchical abstraction, SeSE quantifies the uncertainty inherent in the semantic space after optimal compression and is an expressive generalization of semantic entropy, and provide interpretable, granular uncertainty estimation for LLM long-form generation. Extensive experiments across 24 model-dataset combinations show that SeSE outperforms baseline methods. These findings highlight the potential of SeSE to be a reliable framework for assessing the trustworthiness of LLMs. Future work may explore extending SeSE to multi-agent systems and optimizing the computational overhead.
Appendix A Theoretical Proofs
In this section, we provide a detailed proof of Theorem 1.
Theorem 1 (SeSE Generalizes SE). For any semantic clustering, there exists a semantic graph such that the one-dimensional structural entropy of the graph is equal to the semantic entropy (computed as in Eq. 2111Here, we set logarithms to base 2 for the sake of clarity. Using a different base would only scale entropy values by a constant factor and does not affect relative uncertainty rank.).
Let $C=\{C_{1},\dots,C_{M}\}$ be an arbitrary semantic clustering, and let $p(C_{i}\mid x)$ denote the probability mass assigned to cluster $C_{i}$ (estimated by the frequency of samples falling into each cluster), where $p(C_{i}\mid x)>0$ and $\sum_{i=1}^{M}p(C_{i}\mid x)=1$.
Proof.
Step 1: Construct a semantic graph whose stationary distribution matches $p(C_{i}\mid x)_{i=1}^{M}$. We first construct a directed graph $G_{\text{cluster}}$ where each vertex set represents a semantic cluster $C_{i}\in\{C_{1},\dots,C_{M}\}$. Define a Markov transition matrix $P\in\mathbb{R}^{M\times M}$ by
$P(i,j)=p(C_{j}\mid x),\quad\forall\,i,j\in\{1,\dots,M\}.$
Since $\sum_{j=1}^{M}P(i,j)=\sum_{j=1}^{M}p(C_{j}\mid x)=1$, $P$ is row-stochastic. Consider a distribution $\pi\in\mathbb{R}^{M}$ given by $\pi(i)=p(C_{i}\mid x)$. For any $j$,
$(\pi P)(j)=\sum_{i=1}^{M}\pi(i)P(i,j)=\sum_{i=1}^{M}p(C_{i}\mid x)p(C_{j}\mid x)=p(C_{j}\mid x)\underbrace{\sum_{i=1}^{M}p(C_{i}\mid x)}_{=1}=p(C_{j}\mid x)=\pi(j).$
Hence $\pi$ is a stationary distribution of $P$. Moreover, since $P(i,j)=p(C_{j}\mid x)>0$ for all $i,j$, the chain is irreducible and aperiodic, so the stationary distribution $\pi$ is unique.
Step 2: Apply SeSE with a one-dimensional encoding tree. Consider the one-dimensional encoding tree $\mathcal{T}^{1}$ for $G_{\text{cluster}}$, where the root $\lambda$ contains all clusters, and each cluster $C_{i}$ is represented by a distinct leaf node $\alpha_{i}$ that is directly connected to the root, such that $\mathcal{T}^{1}_{\alpha_{i}}=\{C_{i}\}$. By definition in the main text (Eq. 9), the structural entropy of this single-layer encoding tree $\mathcal{T}^{1}$ is equivalent to the Shannon entropy of the stationary distribution of $G_{\text{cluster}}$:
$H^{\mathcal{T}^{1}}(G_{\text{cluster}})=-\sum_{i=1}^{M}\pi(i)\log_{2}\pi(i).$
When the hierarchy height is restricted to $K=1$, the optimal encoding tree $\mathcal{T}^{\star}$ is uniquely determined as $\mathcal{T}^{1}$. Then,
$\mathrm{SeSE}(G_{\text{cluster}},K=1)=H^{\mathcal{T}^{\star}}(G_{\text{cluster}})=H^{\mathcal{T}^{1}}(G_{\text{cluster}})=-\sum_{i=1}^{M}\pi(i)\log_{2}\pi(i).$
Substituting $\pi(i)=p(C_{i}\mid x)$, we obtain:
$\mathrm{SeSE}(G_{\text{cluster}},K=1)=-\sum_{i=1}^{M}p(C_{i}\mid x)\log_{2}p(C_{i}\mid x)=\mathrm{SE}(x).$
Thus, we have shown that for any semantic clustering $C=\{C_{1},\dots,C_{M}\}$ with distribution $\{p(C_{i}\mid x)\}_{i=1}^{M}$, there exists a corresponding semantic graph such that SeSE with $K=1$ equals SE. ∎
Theorem 1 shows that SeSE can recover SE when $K=1$. Notably, in our implementation, we start by building a response-level semantic graph (i.e, treating each singleton response as a cluster), whose transition probabilities are induced by NLI-based entailment scores between generated responses. This construction captures additional asymmetric and richer semantic dependencies beyond the coarse-grained cluster masses $\{p(C_{i}\mid x)\}$ used by SE. When $K>1$, SeSE can capture the hierarchical structure of the semantic space through a multi-level encoding tree. The resulting optimal tree $\mathcal{T}^{*}$ reflects a progressive semantic partitioning: lower layers correspond to fine-grained partitions that capture subtle distinctions, while higher layers represent broader aggregations that reveal global semantic structural patterns. This hierarchical structure captures both local and global semantic relationships, enhancing the ability to distinguish subtle uncertainty differences. Consequently, SeSE can effectively distinguish uncertainties in complex scenarios where existing methods Farquhar et al. [2024], Nikitin et al. [2024], Li et al. [2025b], Qiu and Miikkulainen [2024] fail to differentiate between superficially similar semantic distributions.
Appendix B Details of SeSE in long-form generation
Figure 5: Overview of SeSE in long-form generation. We decompose the generated long-form response into atomic claims. SeSE considers more sophisticated semantic structural information from the perspective of random walk, and hallucinations are indicated by high SeSE associated with that claim in the constructed bipartite response-claim graph.
In practice, LLMs often output long-form paragraphs containing multiple claims Min et al. [2023]: the smallest semantically distinct unit of information presented within the generations. In long-form generation, we therefore assess uncertainty at the finer-grained claim level rather than simply assigning a single uncertainty score to an entire response or sentence. We have the following observation: given a context $x$, a set of randomly sampled responses $R$, and a set of claims $C$ extracted from the greedily decoded response $r_{T=0}$, we can construct a bipartite graph $G_{cr}=((R,C),E)$ where the edge set $E$ represents semantic entailment relationships between $R$ and $C$. The graph $G_{cr}$ thus captures semantic dependencies between responses $R$ and claims $C$, from which we can extract information that reflects the uncertainty associated with each claim $c\in C$. Intuitively, a claim that is densely connected to the response subgraph (i.e., consistently supported across sampled outputs) is more likely to be factually correct. Conversely, a sparsely connected claim is more likely to be a hallucination. As shown in Fig. 5, we first construct $G_{cr}=((R,C),E)$ and then estimate claim-level uncertainty using SeSE on this graph for hallucination detection.
Response Sampling and Claim Decomposition
Using the same sampling settings as in Section 3.1, we prompt LLMs to sample a stochastic response set $R(\cdot\mid x)$ and a greedily decoded response $r_{T=0}$. Then, we prompt GPT-5-mini with a specific template to decompose $r_{T=0}$ into atomic claims, resulting in the claim set $C$. We adapt the prompt from Jiang et al. [2024]. The prompt is as follows:
    The input is:
    {greedily decoded response}
    Output:
Bipartite Graph Construction
The bipartite graph $G_{cr}$ is constructed by establishing connections between the response set $R$ and the claim set $C$. An edge $e\in E$ is created between a response $r\in R$ and a claim $c\in C$ if $r$ entails $c$. Edge weights are binary: $1$ if entailment holds and 0 otherwise. We leverage the nuanced logical understanding of GPT-5-mini to assess entailment, thus avoiding the brittleness of manually tuned thresholds based on embedding distance or density-based clustering. Specifically, we adapt the following prompt from Manakul et al. [2023] to each pair of response $r\in{R}$ and claim $c\in{C}$ to construct the edge set E{E} of the bipartite graph.
  • Context: {random sampling response}
    Claim: {claim}
    Is the claim supported by the context above?
    Answer Yes or No:
    Output:
And we presents the related ablation results of the entailment estimator in Appendix D.6.
Claim-level Uncertainty Estimation
In the bipartite graph $G_{cr}=((R,C),E)$, we model entailment relations as random walks between response and claim vertices, and quantify the uncertainty of these interactions using structural entropy. By minimizing the $K$-dimensional structural entropy of $G_{cr}$, we obtain its optimal encoding tree $\mathcal{T}^{*}_{cr}$, which captures the inherent hierarchical community structure over $C$ and $R$. Following the same process in Section 3.1, we start by initializing a single-layer encoding tree $\mathcal{T}_{cr}$ in which each leaf node $\gamma$ has the tree root $\lambda$ as its parent. Then, we obtain the optimal $K$-dimensional encoding tree $\mathcal{T}^{*}_{cr}$ using Algorithm 1. In $\mathcal{T}^{*}_{cr}$, the root node $\lambda$ corresponds to the union of claim and response sets, $\mathcal{T}_{\lambda}=R\cup C$. Each leaf node $\gamma$ is a singleton containing an individual claim or response, and intermediate nodes represent hierarchical abstractions at different levels.
The structural entropy associated with each non-root node $\alpha$ quantifies the uncertainty of a random walk transitioning from the parent community $\mathcal{T}_{\alpha^{-}}$ to its child community $\mathcal{T}_{\alpha}$. For any claim $c\in C$, the uncertainty of reaching $c$ is determined by the cumulative entropy of all nodes $\alpha$ encountered along the path from the root node $\lambda$ to the leaf node $\gamma$ with $V_{\gamma}=\{c\}$. Accordingly, we define the SeSE of each claim $c$ as its uncertainty of engaging in random interactions within $G_{cr}$ as detailed below:
$\text{SeSE}\left(G_{cr};c\right)=-\sum_{\alpha\in\mathcal{P}(\lambda\to\gamma)\setminus\{\lambda\}}\frac{g_{\alpha}}{\operatorname{vol}(G_{cr})}\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (17)
Appendix C Details of Hierarchical Abstraction
Merging and Combining Operators Li [2024]
Figure 6: Illustration of the hierarchical abstraction construction with merging and combining operators.